Circular No. 6283

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA CASSIOPEIAE 1995 J. J. Johnson, K. S. Bjorkman, and B. L. Babler, University of Wis-

consin, Madison, report: “CCD spectropolarimetry (range 320–1050 nm, resolution 0.5 nm) of N Cas 1995 (cf. IAUC 6213) was obtained on Dec. 16 and 22 with the University of Wisconsin’s PBO 0.91-m telescope. The spectra obtained on Dec. 16 show strong Hα and Hβ emission, as well as a very strong Balmer absorption jump. In addition, there is a depression in the flux near 430 nm that is believed to be due to overlapping absorption by iron lines. The Ca II infrared triplet is strongly in emission, while Ca H and K are strongly in absorption. The Paschen series is in absorption, except for Paδ, which possibly shows a P-Cyg profile. The polarization is constant with wavelength at 1.5 percent but has changed from previous measurements of N Cas 1995 obtained at PBO, indicating an intrinsic component to the observed polarization. The observations on Dec. 22 were obtained in the midst of the rapid fading of N Cas 1995. The spectra show Hα through Hδ in emission, and the Paschen lines from Paδ upwards are likewise in emission. The Na D absorption line is partially filled in with emission from He I at 587.6 nm. The Balmer absorption jump is very shallow, and the depression due to Fe II at 430 nm seen in the Dec. 16 spectrum is no longer present. The continuum has flattened compared to the Dec. 16 observation. The nova appears to be entering the transition to the nebular phase of its evolution. No lines of O I are seen in either spectrum.”

Visual magnitude estimates (cf. IAUC 6275): 1995 Dec. 19.18 UT, 7.0 (A. Hale, Cloudcroft, NM); 19.75, 7.1 (L. Kiss, Szeged, Hungary); 20.11, 7.1 (R. Fidrich, Budapest, Hungary); 20.73, 7.4 (S. Baroni, Milan, Italy); 21.70 8.4 (Fidrich); 23.12, 8.1 (C. E. Spratt, Victoria, BC); 24.78, 8.4 (J. Van Der Looy, Balen, Belgium); 26.92, 8.7 (M. V. Zanotta, Laino, Italy); 28.09, 9.1 (Fidrich); 29.84, 8.9 (E. Broens, Mol, Belgium); 1996 Jan. 1.78, 8.8 (R. Goncalves, Oeiras, Portugal).

COMET 45P/HONDA-MRKOS-PAJDUŠÁKOVÁ Total visual magnitude estimates: 1995 Dec. 15.06 UT, 8.7 (A. Hale,

Cloudcroft, NM, 10×50 binoculars); 16.10, 8.4 (C. S. Morris, Whitaker Peak, CA, 0.26-m reflector); 21.05, 7.6 (R. Keen, Mt. Thorodin, CO, 0.15-m reflector; 3′ coma); 24.10, 6.6 (C. E. Spratt, Victoria, BC, 0.20-m reflector); 26.08, 6.4 (Spratt, 0.10-m refractor); 28.09, 6.8 (Morris, 20×80 binoculars).

1996 January 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green